A toothpick manufacturer wants every box to contain exactly (on average) 500 toothpicks. In a random sample of 25 boxes, mean is 498 toothpicks and standard deviation is 9 toothpicks.


1
Expert's answer
2021-07-13T05:05:54-0400

One sample t-test.

"H_0:\mu=500."

"H_a:\mu\ne500."

Test statistic: "t=\frac{498-500}{\frac{9}{\sqrt{25}}}=-1.11."

Degrees of freedom: "df=25-1=24."

P-value: "p=2P(T<-1.11)=0.2780."

Since the p-value is greater than 0.05, fail to reject the null hypothesis.

There is no sufficient evidence that the mean number of toothpicks is different of 500.